Currently, you can exchange $1 for either ¥107.20 or €.7622 in New York. In Tokyo, the exchange rate is ¥/€.0074. If you have $1,750, how much profit can you earn with triangle arbitrage?
Answer #1

Steps for Triangular arbitrage are:

Convert Dollar into Yen and get 1,750*107.20 = Yen 187,600

Convert Yen into Euro and get 187,600*0.0074 = Euro 1,388.24

Convert Euro back into Dollar 1,388.24/0.7622 = $1,821.36

Arbitrage profit = 1,821.36-1,750

= $71.36